### v1

@@ Abstract
In this study, we introduce Diamond, a novel method for trustworthy feature interaction discovery. Diamond uniquely integrates the model-X knockoffs framework to control the false discovery rate (FDR), ensuring that the proportion of falsely discovered interactions remains low. We further address the challenges of using off-the-shelf interaction importance measures by proposing a calibration procedure that refines these measures to maintain the desired FDR. Diamond’s applicability spans a wide range of ML models, including deep neural networks, tree-based models, and factorization-based models. Our empirical evaluations on both simulated and real datasets across various biomedical studies demonstrate Diamond’s utility in enabling more reliable data-driven scientific discoveries. This method represents a significant step forward in the deployment of ML models for scientific innovation and hypothesis generation.
@@ 1 Introduction
Recognizing the limitations of univariate interpretations, efforts have been made to extend interpretable ML methods to discover interactions among features. These methods attribute the prediction influence to feature pairs and then rank candidate feature pairs from a trained ML model, with highly ranked pairs indicating higher importance [52, 53, 11, 37, 22, 51, 7, 26, 58]. However, it is important to note that these approaches characterize feature pairs where both features are individually important for a model’s prediction, rather than capturing the synergistic or interactive effects between the two features [54] (see the detailed explanation in Sec.4.2). Furthermore, the induced ranked list of feature pairs must be cut off at a certain confidence level for use in scientific discovery and hypothesis validation [52]. However, selecting this threshold is typically under user control and must be set arbitrarily. Worse still, existing methods are sensitive to perturbations, in the sense that even imperceivable, random perturbations of the input data may lead to dramatic changes in the importance ranking [19, 25, 35].
@@
Three components of Diamond are novel (Fig. 1). First, Diamond achieves FDR control by leveraging the model-X knockoffs framework [2, 6]. The core idea of this framework is to generate dummy features that perfectly mimic the empirical dependence structure among the original features while being conditionally independent of the response given the original features. Second, we discover that naively using off-the-shelf feature interaction importance measures cannot correctly control the FDR. To address this issue, we propose a calibration procedure to distill non-additive or interactive effects from the reported interaction importance measures from existing methods, thereby maintaining FDR control at the target level. Third, Diamond is applicable to a wide range of ML models, including deep neural networks (DNNs), tree-based models, and factorization-based models. We have applied Diamond to various simulated and real datasets to demonstrate its empirical utility. Practically speaking, Diamond paves the way for the wider deployment of machine learning models in scientific discovery and hypothesis generation.
Figure 1: Overview of Diamond. (a) Diamond achieves false discovery rate (FDR) control by leveraging knockoffs – dummy features that perfectly mimic the empirical dependence structure among the original features but are conditionally independent of the response given the original features. (b) Diamond trains generic ML models using both the original features and their knockoff counterparts as inputs. Diamond quantifies feature interactions from trained ML models and produces a ranked list of these interactions with estimated FDR, allowing users to confidently determine a cutoff threshold based on their desired confidence level. (c) Diamond uses a calibration procedure to distill non-additive or interactive effects from the reported interaction importance measures from existing methods, thereby maintaining FDR control at the target level. The calibration process is designed to remove both prediction-dependent marginal effects and prediction-independent feature biases from the reported feature interactions, leaving only the prediction-dependent non-additive interactive effects.
@@ 2.1 Diamond overview
In this study, we focus on interpreting the fitted ML model for data-driven scientific discovery by identifying the non-additive feature interactions that contribute most to the model’s predictions. We say that $\mathcal{I}\subset\{1,\cdots,p\}$ is a non-additive interaction of function $f$ if and only if $f$ cannot be decomposed into an addition of $\left|\mathcal{I}\right|$ subfunctions $f_{i}$, each of which excludes a corresponding interaction feature [49, 52], i.e., $f(x)\neq\sum_{i\in\mathcal{I}}f_{i}\left(x_{\{1,\cdots,p\}\backslash i}\right)$. For example, the multiplication between two features $x_{i}$ and $x_{j}$ is a non-additive interaction because it cannot be decomposed into a sum of univariate functions, i.e., $x_{i}x_{j}\neq f_{i}(x_{j})+f_{j}(x_{i})$. Assume that there exists a group of interactions $\mathcal{S}=\left\{\mathcal{I}_{1},\mathcal{I}_{2},\cdots\right\}$ such that conditional on interactions $\mathcal{S}$, the response $\mathbf{Y}$ is independent of interactions in the complement $\mathcal{S}^{c}=\{1,\cdots,p\}\times\{1,\cdots,p\}\backslash\mathcal{S}$. In this setting, our goal is to accurately discover feature interactions in $\mathcal{S}$ without erroneously reporting too many incorrect interactions in $\mathcal{S}^{c}$.
@@
Diamond trains a generic ML model that takes as input an augmented data matrix $(\mathbf{X},\tilde{\mathbf{X}})\in\mathbb{R}^{n\times 2p}$, created by concatenating the original data matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ with its knockoffs $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times p}$. After training, Diamond quantifies feature interactions by interpreting the trained model and produces a ranked list of these interactions. Diamond subsequently uses a calibration procedure to distill non-additive or interactive effects from the reported interaction importance measures from existing methods, thereby maintaining FDR control at the target level (Fig. 1b). The calibration process is designed to remove both prediction-dependent marginal effects and prediction-independent feature biases from the reported feature interactions, leaving only the prediction-dependent non-additive interactive effects (Fig. 1c).
@@ 2.2 Diamond discovers FDR-controlled interactions on simulated datasets
Figure 2: Evaluating Diamond for FDR control on simulated datasets. (a) The evaluation is based on a test suite of 10 data-generating simulation functions proposed by [52]. (b) Diamond identifies important non-additive interactions with controlled FDR, compatible with various ML models. The calibration procedure is critical; without it, the FDR cannot be controlled. (c) The reported interaction importance from existing methods in simulation function $F_{1}$ reveals a clear distribution disparity between original-only interactions and those involving knockoffs. The proposed calibration procedure mitigates the disparity by distilling non-additive interactive effects from the reported interaction importance measures. (d) Baseline methods fail to correctly control the FDR, thereby rendering the reported high power and AUROC invalid.
@@
Our analysis shows that Diamond consistently identifies important non-additive interactions with controlled FDR across various ML models (Fig. 2b). Under controlled FDR, the multi-layer perceptron (MLP) model demonstrates better performance in identifying important interactions than other ML models, as measured by the statistical power and the area under the receiver operating characteristic curve (AUROC). This superior performance is attributed to the inclusion of a knockoff-tailored plugin pairwise-coupling layer design [33], which has been shown to maximize statistical power (See details in Sec. S.1). We discovered that the proposed calibration procedure (Sec 4.2) is critical; without it, the FDR cannot be controlled by naively using reported interaction importance values from existing methods (Fig. 2b and Fig. S.2).
To gain insight into the FDR control failure in the absence of calibration, we conducted a qualitative comparison assessing interaction importance before and after calibration using the simulation function $F_{1}$ (Fig. 2c and Fig. S.3). The primary cause of the FDR control failure appears to lie in the distributional disparity between original-only interactions (i.e., interactions involving two features from the original feature set rather than knockoffs) and interactions involving knockoffs. This observation suggests a violation of the knockoff filter’s assumption in controlling the FDR (See discussion in Sec 4.2). The proposed calibration procedure mitigates the disparity by extracting non-additive interactive effects from the reported interaction importance measures, thereby enhancing the utility of knockoff-involving interactions as a negative control for FDR estimation.
Finally, we verified whether alternative baseline methods can accurately identify important non-additive interactions with controlled FDR. We compared two baseline methods for FDR estimation: one based on permutation-based interaction-wise p-values coupled with the Benjamini-Hochberg procedure, and the other representing interaction-wise FDR as the aggregation of feature-wise FDR (See details in Sec. 4.6). Our analysis shows that neither method correctly controls the FDR. This greatly reduces the utility of these methods, despite their reported high power and AUROC.
@@ 2.3 Diamond is robust to various knockoffs and importance measures
Figure 3: Evaluating Diamond for robustness on simulated datasets. (a) Diamond is applied with an MLP model using knockoffs generated by various approaches, including KnockoffsDiagnostics, KnockoffGAN, Deep knockoffs, and VAE knockoffs. Diamond consistently discovers important non-additive interactions with controlled FDR and comparable statistical power across various knockoff designs. (b) Diamond is applied with an MLP model and interpreted using three methodologically different interaction importance measures: Expected Hessian, Integrated Hessian, and model-specific measures. Diamond consistently discovers important non-additive interactions with controlled FDR and comparable statistical power across various interaction importance measures.
@@
The design of Diamond incorporates several key components, including knockoff generation and interaction importance measures. To demonstrate the robustness of Diamond in FDR control, we conducted a control study where we modified Diamond by replacing these two components with alternative solutions. Specifically, we considered two variants of Diamond using the MLP model. In the first setting, we replaced KnockoffsDiagnostics with three alternatives: KnockoffGAN [23], Deep knockoffs [42] or VAE knockoffs [31], while keeping the MLP model and the interaction importance measure (i.e., Expected Hessian [14]) unchanged. In the second setting, we replaced the interaction importance measure, Expected Hessian, with two alternatives: Integrated Hessian [22] and a model-specific interaction importance measure derived from the MLP weights (See details in Sec. S.2), while keeping the MLP model and the knockoff generation unchanged. For each variant, we applied Diamond to the test suite of 10 simulated datasets using the same settings.
@@
The results indicate that Diamond is robust to various knockoff designs and importance measures. In the first study, our analysis shows that Diamond consistently discovers important non-additive interactions with controlled FDR across various knockoff designs (Fig. 3a). Despite the methodological differences in knockoff generation between KnockoffsDiagnostics, KnockoffGAN, Deep knockoffs, and VAE knockoffs, Diamond achieves comparable statistical power and AUROC in each case. This result demonstrates the Diamond’s stability in identifying important interactions in knockoff-based FDR estimation.
@@
In the second study, our analysis shows that Diamond discovers important non-additive interactions with controlled FDR across various interaction importance measures (Fig. 3b). Despite the methodological differences between Expected Hessian, Integrated Hessian, and model-specific measures, which involve mechanistic disparities in elucidating the relationships between features and responses, Diamond achieves valid FDR control while maintaining reasonably good statistical power and AUROC. This result demonstrates Diamond’s flexibility in identifying important interactions based on interaction-wise importance ranking.
@@ 2.4 Diamond identifies interactions to explain disease progression
Figure 4: Evaluating Diamond on a real diabetes dataset. (a) Diamond is applied with four different ML models to identify important non-additive interactions and is compared against two baseline methods. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top interaction annotated. The blue stars indicate interactions supported by literature evidence, referenced by the accompanying PubMed identifiers. (b) Each feature contributes differently to predicting disease progression, as measured by the Expected Gradient scores in the MLP model. It is worth mentioning that two marginally important features do not necessarily result in important interactions, as anticipated in Diamond’s design. (c) The top interaction, between body mass index and serum triglycerides level, is qualitatively evaluated from four aspects: the contribution of feature values to response prediction, the marginal importance measure, the interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.
@@
We then evaluated the performance of Diamond on a real dataset [13], assessing its ability to identify important interactions in the progression of disease. We used a quantitative study with $n=442$ diabetes patients, each characterized by $10$ standardized baseline features, including age, sex, body mass index, average blood pressure, and six blood serum measurements (Fig. 4b). The task is to construct an ML model to predict the response of interest, which is a quantitative measure of disease progression one year after the baseline. We considered four different types of ML models: MLP, LightGBM, XGBoost, and factorization machines (FM). For robustness, we repeated each experiment 20 times using different random seeds. Each repetition involved knockoff generation, ML model training, and interaction-wise FDR estimation.
@@
Our analysis shows that Diamond consistently identifies the important interaction between body mass index (BMI) and serum triglycerides level (STL) across three of the four different ML models (Fig. 4a). The only exception is the FM method, for which Diamond fails to identify any important interactions. We investigated the identified interaction via literature search and found that high BMI and high STL are associated with an increased risk of diabetes, as reported by multiple studies [17, 59, 20]. The literature evidence is further supported by the qualitative evaluation (Fig. 4c). Specifically, we examined the identified interaction in four ways: the contribution of feature values to response prediction, the marginal importance measure, the interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure. This analysis showed that high BMI and high STL contribute to more severe progression of diabetes. Meanwhile, higher STL further reinforces the contribution of BMI to the progression of diabetes synergistically.
@@
Finally, we find that Diamond does not necessarily report a lower FDR for interactions composed of two marginally important features. All $10$ baseline features from diabetes patients are effective in predicting disease progression, as measured by the Expected Gradient scores in the MLP model (Fig. 4b). Among these features, in addition to BMI and STL, blood pressure, blood sugar level, and high-density lipoproteins are more predictive than others. It is important to note that Diamond still associates most interactions composed of these highly predictive features with a high FDR threshold.

### v2

@@ Abstract
In this study, we introduce Diamond, a novel method for trustworthy feature interaction discovery. Diamond uniquely integrates the model-X knockoffs framework to control the false discovery rate (FDR), ensuring that the proportion of falsely discovered interactions remains low. A key innovation in Diamond is its non-additivity distillation procedure, which refines existing interaction importance measures to distill non-additive interaction effects, ensuring that FDR control is maintained. This approach addresses the limitations of off-the-shelf interaction measures, which, when used naively, can lead to inaccurate discoveries. Diamond’s applicability spans a wide range of ML models, including deep neural networks, transformer models, tree-based models, and factorization-based models. Our empirical evaluations on both simulated and real datasets across various biomedical studies demonstrate Diamond’s utility in enabling more reliable data-driven scientific discoveries. This method represents a significant step forward in the deployment of ML models for scientific innovation and hypothesis generation.
@@ 1 Introduction
Recognizing the limitations of univariate interpretations, efforts have been made to extend interpretable ML methods to discover interactions among features. These methods attribute the prediction influence to feature pairs and then rank candidate feature pairs from a trained ML model, with highly ranked pairs indicating higher importance [59, 60, 13, 42, 25, 58, 9, 30, 65]. However, it is important to note that these approaches characterize feature pairs where both features are individually important for a model’s prediction, rather than capturing the synergistic or interaction effects between the two features [61] (see the detailed explanation in Sec.4.2). Furthermore, the induced ranked list of feature pairs must be cut off at a certain confidence level for use in scientific discovery and hypothesis validation [59]. However, selecting this threshold is typically under user control and must be set arbitrarily. Worse still, existing methods are sensitive to perturbations, in the sense that even imperceivable perturbations of the input data may lead to dramatic changes in the importance ranking [21, 29, 40].
@@
Three components of Diamond are novel (Fig. 1). First, Diamond achieves FDR control by leveraging the model-X knockoffs framework [3, 8]. The core idea of this framework is to generate dummy features that perfectly mimic the empirical dependence structure among the original features while being conditionally independent of the response given the original features. Second, we discover that naively using off-the-shelf feature interaction importance measures cannot correctly control the FDR.
To address this issue, we distill non-additive interaction effects from the interaction importance measures reported by existing methods, ensuring FDR control at the target level. Third, Diamond is applicable to a wide range of ML models, including deep neural networks (DNNs), Kolmogorov-Arnold networks (KANs) [36], transformer models [22], tree-based models, and factorization-based models.
We have applied Diamond to various simulated and real datasets to demonstrate its empirical utility. Practically speaking, Diamond paves the way for the wider deployment of machine learning models in scientific discovery and hypothesis generation.
Figure 1: Overview of Diamond. (a) Diamond achieves false discovery rate (FDR) control by leveraging knockoffs – dummy features that perfectly mimic the empirical dependence structure among the original features but are conditionally independent of the response given the original features. (b) Diamond trains generic ML models using both the original features and their knockoff counterparts as inputs. Diamond quantifies feature interactions from trained ML models and produces a ranked list of these interactions with estimated FDR, allowing users to confidently determine a cutoff threshold based on their desired confidence level. (c) Existing feature importance measures are unable to directly capture non-additive interactions. Geometrically, the marginal or interaction importance reported by current interpretation methods correspond to the projection of the total contribution to the prediction onto a one-dimensional feature axis or a two-dimensional feature-feature plane. Non-additive effects, however, represent the difference between the interaction effects and the marginal effects. These non-additive effects can manifest as either synergistic or repressive interactions. (d) Diamond distills non-additive effects from the interaction importance measures reported by existing methods, thereby maintaining FDR control at the target level. The distillation procedure is designed to remove both label-dependent marginal effects and label-independent feature biases from the reported feature interactions, leaving only the label-dependent non-additive interaction effects.
@@ 2.1 Diamond overview
In this study, we focus on interpreting the fitted ML model for data-driven scientific discovery by identifying the non-additive feature interactions that contribute most to the model’s predictions. We say that $\mathcal{I}\subset\{1,\cdots,p\}$ is a non-additive interaction of function $f$ if and only if $f$ cannot be decomposed into an addition of $\left|\mathcal{I}\right|$ subfunctions $f_{i}$, each of which excludes a corresponding interaction feature [56, 59], i.e., $f(x)\neq\sum_{i\in\mathcal{I}}f_{i}\left(x_{\{1,\cdots,p\}\backslash i}\right)$. For example, the multiplication between two features $x_{i}$ and $x_{j}$ is a non-additive interaction because it cannot be decomposed into a sum of univariate functions, i.e., $f(x_{i},x_{j})=x_{i}x_{j}\neq f_{i}(x_{j})+f_{j}(x_{i})$. On the other hand, $\log(x_{i}x_{j})$ is considered to be an additive interaction because it can be decomposed in the logarithmic space, i.e., $f(x_{i},x_{j})=\log(x_{i}x_{j})=\log(x_{j})+\log(x_{i})$. Assume that there exists a group of interactions $\mathcal{S}=\left\{\mathcal{I}_{1},\mathcal{I}_{2},\cdots\right\}$ such that, conditional on interactions $\mathcal{S}$, the response $\mathbf{Y}$ is independent of interactions in the complement $\mathcal{S}^{c}=\{1,\cdots,p\}\times\{1,\cdots,p\}\backslash\mathcal{S}$. In this setting, our goal is to accurately discover feature interactions in $\mathcal{S}$ without erroneously reporting too many incorrect interactions in $\mathcal{S}^{c}$.
@@
Diamond trains a generic ML model that takes as input an augmented data matrix $(\mathbf{X},\tilde{\mathbf{X}})\in\mathbb{R}^{n\times 2p}$, created by concatenating the data matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ with its knockoffs $\tilde{\mathbf{X}}\in\mathbb{R}^{n\times p}$ along the feature axis. After training, Diamond quantifies feature interactions by interpreting the trained model and produces a ranked list of these interactions (Fig. 1b). However, existing feature importance measures cannot directly capture non-additive interactions. Geometrically, the marginal or interaction importance reported by current interpretation methods correspond to the projection of the total contribution to the prediction onto a one-dimensional feature axis or a two-dimensional feature-feature plane. Non-additive effects, however, represent the difference between the interaction effects and the marginal effects (Fig. 1c). Hence, Diamond distills non-additive interaction effects from the reported interaction importance measures from existing methods, thereby maintaining FDR control at the target level. The distillation process is designed to remove both label-dependent marginal effects and label-independent feature biases from the reported feature interactions, leaving only the label-dependent non-additive interaction effects (Fig. 1d).
@@ 2.2 Diamond discovers FDR-controlled interactions on simulated datasets
Figure 2: Evaluating Diamond for FDR control on simulated datasets. (a) The evaluation is based on a test suite of 10 data-generating simulation functions proposed by [59]. (b) The reported interaction importance from existing methods in simulation function $F_{1}$ reveals a clear distribution disparity between original-only interactions and those involving knockoffs. The distilled non-additive interactions help mitigate distributional disparities. (c) Baseline methods fail to correctly control the FDR, thereby rendering the reported high power and AUROC invalid. (d) Diamond identifies important non-additive interactions with controlled FDR, compatible with various ML models. The non-additivity distillation procedure is critical; without it, the FDR cannot be controlled. Even in scenarios of model misspecification, such as using convolutional neural networks for tabular data, Diamond maintains FDR control, albeit with a loss of power.
@@
Our analysis shows that Diamond consistently identifies important non-additive interactions with controlled FDR across all ML models (Fig. 2d). Under controlled FDR, the multi-layer perceptron (MLP) and Feature Tokenizer Transformer (FT-Transformer) models [22] demonstrate better performance in identifying important interactions than other ML models, as measured by the statistical power and the area under the receiver operating characteristic curve (AUROC). This superior performance is attributed to the inclusion of a knockoff-tailored plugin pairwise-coupling layer design [38], which has been shown to maximize statistical power (See details in Sec. A.1). It is important to note that certain ML models, such as convolutional neural networks, are not well-suited for modeling tabular data due to their design for capturing local patterns in images or text. Even in the presence of model misspecification, Diamond maintains FDR control, albeit with a loss of power. This highlights Diamond’s robustness and its broad applicability across various ML models.
We discovered that the proposed non-additivity distillation (Sec 4.2) is critical; without it, the FDR cannot be controlled by naively using reported interaction importance values from existing methods (Fig. 2d and Fig. A.2). To gain insight into the FDR control failure in the absence of non-additivity distillation, we conducted a qualitative comparison assessing interaction importance before and after distillation using the simulation function $F_{1}$ (Fig. 2b and Fig. A.3). The primary cause of the FDR control failure appears to lie in the distributional disparity between original-only interactions (i.e., interactions involving two features from the original feature set rather than knockoffs) and interactions involving knockoffs. This observation suggests a violation of the knockoff filter’s assumption in controlling the FDR (See discussion in Sec 4.2). The proposed non-additivity distillation procedure mitigates the disparity by extracting non-additive interaction effects from the reported interaction importance measures, thereby enhancing the utility of knockoff-involving interactions as a negative control for FDR estimation.
Finally, we verified whether alternative baseline methods can accurately identify important non-additive interactions with controlled FDR (Fig. 2c). We compared three baseline methods for FDR estimation: two relying on permutation-based interaction-wise p-values, combined with the Benjamini-Hochberg and the Benjamini-Yekutieli procedures respectively, and one that represents interaction-wise FDR as an aggregation of feature-wise FDR (See details in Sec. 4.7). Our analysis reveals that none of these baseline methods effectively control the FDR. This greatly reduces the utility of these methods, despite their reported high power and AUROC.
@@ 2.3 Diamond is robust to various knockoffs and importance measures
Figure 3: Evaluating Diamond for robustness on simulated datasets. (a) Diamond demonstrates robustness across knockoffs generated by various methods, including KnockoffsDiagnostics, KnockoffGAN, Deep knockoffs, and VAE knockoffs. (b) Diamond maintains FDR control when paired with invalid knockoffs, generated by independently permuting each feature across samples, albeit at the cost of reduced power. (c) Diamond demonstrates robustness across methodologically different interaction importance measures: Expected Hessian, Integrated Hessian, and model-specific measures.
@@
The design of Diamond incorporates several key components, including knockoff generation and interaction importance measures. To demonstrate the robustness of Diamond in FDR control, we conducted a control study where we modified Diamond by replacing these two components with alternative solutions. Specifically, we considered two variants of Diamond using the MLP model. In the first setting, we replaced KnockoffsDiagnostics with three alternatives: KnockoffGAN [26], Deep knockoffs [48] or VAE knockoffs [35], while keeping the MLP model and the interaction importance measure (i.e., Expected Hessian [16]) unchanged. In the second setting, we replaced the interaction importance measure, Expected Hessian, with two alternatives: Integrated Hessian [25] and a model-specific interaction importance measure derived from the MLP weights (See details in Sec. A.2), while keeping the MLP model and the knockoff generation unchanged. For each variant, we applied Diamond to the test suite of 10 simulated datasets using the same settings.
@@
The results indicate that Diamond is robust to various knockoff designs and importance measures. In the first study, our analysis shows that Diamond consistently discovers important non-additive interactions with controlled FDR across various knockoff designs (Fig. 3a). Despite the methodological differences in knockoff generation between KnockoffsDiagnostics, KnockoffGAN, Deep knockoffs, and VAE knockoffs, Diamond achieves comparable statistical power and AUROC in each case. This result demonstrates the Diamond’s stability in identifying important interactions in knockoff-based FDR estimation. Furthermore, it is worth noteworthy that Diamond is robust even when the knockoffs are poorly generated. Specifically, we generated invalid knockoffs by independently permuting each feature across samples, thereby violating the definition of knockoffs outlined in Sec. 4.1. Our analysis demonstrates that even in this extreme scenario, Diamond maintains FDR control, albeit with reduced power (Fig. 3b).
@@
In the second study, our analysis shows that Diamond robustly identifies FDR-controlled non-additive interactions across a range of interaction importance measures (Fig. 3c). Despite the methodological differences between Expected Hessian, Integrated Hessian, and model-specific measures, which involve mechanistic disparities in elucidating the relationships between features and responses, Diamond achieves valid FDR control while maintaining reasonably good statistical power and AUROC. This result demonstrates Diamond’s flexibility in identifying interactions based on various importance measures.
@@ 2.4 Diamond identifies interactions to explain disease progression
Figure 4: Evaluating Diamond on a real diabetes dataset. (a) Each feature contributes differently to predicting disease progression, as measured by the Expected Gradient scores in the MLP model. (b) Diamond is compared against three baseline methods. The blue stars indicate interactions supported by literature evidence, referenced by the accompanying PubMed identifiers. (c) Diamond is used with various ML models to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top interaction annotated. It is worth mentioning that two marginally important features do not necessarily result in important interactions, as anticipated in Diamond’s design. (d) The top interaction, between body mass index and serum triglycerides level, is qualitatively evaluated from three aspects: the marginal importance measure, the interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure.
@@
We then evaluated the performance of Diamond on a real dataset [15], assessing its ability to identify important interactions in the progression of disease. We used a quantitative study with $n=442$ diabetes patients, each characterized by $10$ standardized baseline features, including age, sex, body mass index, average blood pressure, and six blood serum measurements (Fig. 4a). The task is to construct an ML model to predict the response of interest, which is a quantitative measure of disease progression one year after the baseline. We considered different types of ML models: MLP, FT-Transformer, KAN, LightGBM, XGBoost, and factorization machines (FM). We assessed the trained ML model using 5-fold cross-validation and selected the one with the best performance for model interpretation and interaction discovery. For robustness, we repeated each experiment 20 times using different random seeds. Each repetition involved knockoff generation, ML model training, and interaction-wise FDR estimation.
@@
Our analysis shows that Diamond identifies the important interaction between body mass index (BMI) and serum triglycerides level (STL) across three different ML models (Fig. 4c). We investigated the identified interaction via literature search and found that high BMI and high STL are associated with an increased risk of diabetes, as reported by multiple studies [19, 66, 23]. The literature evidence is further supported by the qualitative evaluation (Fig. 4d). Specifically, we examined the identified interactions from three perspectives: the marginal importance measure, the interaction importance measure, and the contribution of the marginal importance measures to the interaction importance measure. This analysis showed that high BMI and high STL contribute to more severe progression of diabetes. Meanwhile, higher STL further reinforces the contribution of BMI to the progression of diabetes synergistically.
Unlike other ML models, the FT-Transformer identifies two distinct interactions: the interaction between blood pressure and high-density lipoproteins, and the interaction between age and sex. We investigated the identified interactions through a literature search and found that high-density lipoprotein levels are positively associated with hypertension in individuals with elevated levels of circulating CD34-positive cells [53]. We also found that the prevalence of diabetes varies by sex, with men typically being diagnosed at a younger age and lower BMI, while young women with diabetes are currently less likely than men to receive treatment [27].
@@
Finally, we find that Diamond does not necessarily report a lower FDR for interactions composed of two marginally important features. All $10$ baseline features from diabetes patients are effective in predicting disease progression, as measured by the Expected Gradient scores in the MLP model (Fig. 4a). Among these features, in addition to BMI and STL, blood pressure, blood sugar level, and high-density lipoproteins are more predictive than others. It is important to note that Diamond still associates most interactions composed of these highly predictive features with a high FDR threshold.